# DESIGNER: Design-Logic-Guided Multidisciplinary Data Synthesis for LLM Reasoning

### v3

@@ 6 Related Work
Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches. Query-centric approaches iteratively expand a seed query pool: Self-Instruct (Wang et al., 2023) samples from an initial pool to generate new QA pairs, while Wizard LM (Luo et al., 2024; Xu et al., 2023; Luo et al., 2025) and Auto Evol-Instruct (Zeng et al., 2024) evolve instructions for greater diversity and complexity. CoT-Self-Instruct (Yu et al., 2025) integrates reasoning into generation, Prismatic Synthesis (Jung et al., 2025) emphasizes gradient-level diversity, and SPARQ (Havrilla et al., 2025) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance. However, these methods remain constrained by the coverage of the seed pool and the inherent model biases. Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (Ding et al., 2023) generates world-knowledge questions from sources like Wikidata, Humpback (Li et al., 2024) infers instructions backward from documents, MAmmoTH2 (Yue et al., 2024) mines web content via a Recall-Extract-Refine pipeline. NaturalReasoning (Yuan et al., 2025) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template. Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (Huang et al., 2025), and the structurally guided approaches of Xu et al. (2025), Wang et al. (2025), and Bu et al. (2025). However, these approaches have two primary limitations compared to our method. First, they are query-centric methods that synthesize new questions from seed questions, which inherently constrains them to the scope of the initial seed pool and prevents generation from richer raw corpora. Second, their methodologies are tailored for disciplines with strong logical and structural reasoning, such as mathematics and coding, and cannot be readily extended to arbitrary multidisciplinary fields, such as the humanities, social sciences, and medicine. In contrast, our method synthesizes challenging and diverse questions by introducing multidisciplinary design logics to align with the corresponding documents. This approach is discipline-agnostic, making it applicable to any discipline.
@@
Another line of work leverages synthetic long CoT data to distill reasoning abilities into smaller models. DeepSeek (Guo et al., 2025) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (Moshkov et al., 2025) augments math ability by generating long CoT traces for AoPS problems. OmniThought (Cai et al., 2025) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (Guha et al., 2025) systematically explores recipes for generating long reasoning traces across these domains. While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
@@ 7 Conclusion
In this paper, we introduced DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs. Our core innovation is the concept of “Design Logic”, which abstracts the strategic process that human experts use to create challenging questions. By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources. We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning.
@@
We provide all complete and detailed data processing steps used for dataset synthesis in the main paper and the appendix. To facilitate reproducibility and independent verification, we have released a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest. The released data includes our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data.
@@ Appendix C Design Logic and Question Quantity
Number of Design Logics, Book Questions, and Web Questions by Discipline
@@ Appendix H Experimental Details
MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (Wang et al., 2024b) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
Appendix I Prompts
@@
Appendix J Design Logic Examples

### v5

@@ 6 Related Work
Existing data synthesis methods can be broadly grouped into query-centric and doc-centric approaches. Query-centric approaches iteratively expand a seed query pool: Self-Instruct (wang2023self) samples from an initial pool to generate new QA pairs, while Wizard LM (luowizardcoder; xu2023wizardlm; luowizardmath) and Auto Evol-Instruct (zeng2024automatic) evolve instructions for greater diversity and complexity. CoT-Self-Instruct (yu2025cot) integrates reasoning into generation, Prismatic Synthesis (jung2025prismatic) emphasizes gradient-level diversity, and SPARQ (havrilla2025synthetic) evaluates difficulty and diversity, linking them to in- and out-of-distribution performance. However, these methods remain constrained by the coverage of the seed pool and the inherent model biases. Doc-centric approaches instead derive questions directly from raw corpora, ensuring stronger factual grounding: UltraChat (ding2023enhancing) generates world-knowledge questions from sources like Wikidata, Humpback (li2024self) infers instructions backward from documents, MAmmoTH2 (yue2024mammoth2) mines web content via a Recall-Extract-Refine pipeline. NaturalReasoning (yuan2025naturalreasoning) focuses on generating high-difficulty reasoning questions, but its diversity is inherently limited by a single prompt template. Furthermore, several studies have explored generating new questions through external guidance, such as KPDDS (huang2025key), and the structurally guided approaches of xu2025synthesis, wang2025rv, and bu2025enhanced. However, these approaches have two primary limitations compared to our method. First, they are query-centric methods that synthesize new questions from seed questions, which inherently constrains them to the scope of the initial seed pool and prevents generation from richer raw corpora. Second, their methodologies are tailored for disciplines with strong logical and structural reasoning, such as mathematics and coding, and cannot be readily extended to arbitrary multidisciplinary fields, such as the humanities, social sciences, and medicine. In contrast, our method synthesizes challenging and diverse questions by introducing multidisciplinary design logics to align with the corresponding documents. This approach is discipline-agnostic, making it applicable to any discipline.
@@
Another line of work primarily investigates the synthesis of CoT reasoning data. DeepSeek (guo2025deepseek) transfers reasoning skills from DeepSeek-R1 to open-source models, while OpenMathReasoning (moshkov2025aimo) augments math ability by generating long CoT traces for AoPS problems. OmniThought (cai2025reasoning) collects multidisciplinary questions (math, code, science) to build a large CoT distillation dataset, and OpenThoughts (guha2025openthoughts) systematically explores recipes for generating long reasoning traces across these domains. While these approaches primarily focus on generating high-quality reasoning processes for existing questions, they lack the capability to produce original, diverse, and multidisciplinary questions, which is the central focus of our method.
@@ 7 Conclusion
In this paper, we introduced DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs. Our core innovation is the concept of “Design Logic”, which abstracts the strategic process that human experts use to create challenging questions. By leveraging design logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from raw text sources. We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning. Importantly, the value of design logic lies not only in its effect on final performance but also in its controllable, abstractable, and generalizable synthesis mechanism. The guidance of the design logic is crucial for ensuring the difficulty and diversity of the synthesized questions. Section 4.1 shows that our design logics allow precise control over problem difficulty, enabling the generation of highly challenging questions far beyond baseline datasets and commonly used benchmarks. Section 4.2 shows that design-logic-guided synthesis achieves higher scores on all five diversity metrics compared with baseline datasets.
@@
We provide all complete and detailed data processing steps used for data synthesis in the main paper and in the appendix. To facilitate reproducibility and independent verification, we have released a subset of our synthesized data that does not involve legal compliance concerns or potential conflicts of interest. The released data include our synthesized questions, reference answers, long CoT responses, the corresponding source corpora and design logic, as well as associated labels (e.g., discipline, difficulty, question type), enabling reproduction and validation of our method’s effectiveness and the quality of the synthesized data. Furthermore, we have released a design-logic library that contains all design logics extracted in this study, together with discipline, difficulty, and question type labels, enabling the research community to directly match these design logics to their own raw documents and synthesize new questions that exhibit the same complex reasoning.
@@ Appendix C Design Logic and Question Quantity
Number of Design Logics, book questions, and web questions by discipline.
@@ Appendix H Experimental Details
MMLU (hendrycks2020measuring) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (wang2024mmlu) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
GPQA-Diamond (rein2024gpqa) Physics, Chemistry, Biology 10
GPQA-Main (rein2024gpqa) Physics, Chemistry, Biology 10
SuperGPQA (du2025supergpqa) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
Appendix I Performance on GSM8K and MATH-500
We evaluated the models trained on our data in Table 2 on GSM8K and MATH-500.
Table 11: The SFT experiment results on GSM8K and MATH-500. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web + DLR-Book dataset.
Model GSM8K MATH-500
[…]
Llama-3.2-3B-SFT (DLR-Book) 81.05 $5±\pm1.4$2
[…]
GSM8k and MATH-500 are exclusively focused on mathematical problems. Our dataset’s objective is not mathematics; rather, it is designed to create a multidisciplinary dataset. Nevertheless, the model trained on our data still exhibits competitive performance in the mathematical domain.
For MATH500, the slightly lower performance of our models compared with the official Qwen3 models that have undergone the full post-training process is likely due to the fact that the official Qwen3 models were trained with a substantially larger volume of high-quality, expert-written mathematical data and received more extensive post-training. In contrast, mathematics represents only about 10% of our dataset, and our math questions are synthesized from raw web and book corpora rather than curated by human experts, which naturally leads to lower quality relative to professional math-focused training data.
Notably, models trained on either DLR-Book or DLR-Web+Book achieve 95.60% on GSM8K, matching the official Qwen3-8B score. GSM8K performance is already near saturation. Prior work (vendrow2025large) reports that approximately 5% of GSM8K questions are mislabeled or ambiguous, which makes 100% accuracy unattainable and positions 95.60% near the practical upper bound. Even stronger LLMs fall within a similar range. For instance, according to the Llama 3 technical report (dubey2024llama), GPT-4o reaches 96.1 on GSM8K under 8-shot CoT setting.
Appendix J Prompts
@@ Appendix I Performance on GSM8K and MATH-500
Prompt for Question Answerability Evaluation Your task is to review a question and decide whether it provides sufficient conditions to derive solutions. If the question is a complete and answerable one, output ‘Yes’. Open-ended questions are considered answerable. If it is an incorrect question that cannot be answered, output ‘No’. Output only ‘Yes’ or ‘No’.
‘‘question’’: ‘‘{question}’’
Output:
Figure 14: The prompt used to evaluate whether the synthesized questions are complete and answerable. The placeholder {question} is replaced with the specific synthesized question.
Prompt for Design Logic Consistency Evaluation Your task is to judge whether a ‘‘question’’ is designed according to the Mermaid-format ‘‘design logic’’. You should ignore superficial wording changes. If the logic is consistent or if there are implicit references that preserve consistency, output ‘Yes’. If the design logic underlying the question is not consistent with the ‘‘design logic’’, output ‘No’. Output only ‘Yes’ or ‘No’.
‘‘design logic’’: ‘‘{design_logic}’’
‘‘question’’: ‘‘{question}’’
Output:
Figure 15: The prompt used to evaluate the consistency between the synthesized question and its corresponding design logic. The placeholders {design_logic} and {question} are replaced with the specific design logic and the corresponding synthesized question, respectively.
Prompt for Answer Consistency Evaluation You are an expert reviewer who compares a ‘‘reference_answer’’ with a model ‘‘response’’ for the same question. Decide whether both provide the same final answer. Ignore differences in reasoning or wording, and output ‘Yes’ when the final answers or key conclusions are identical. Output ‘No’ when the conclusions differ. Output only ‘Yes’ or ‘No’.
‘‘question’’: ‘‘{question}’’
‘‘reference_answer’’: ‘‘{reference_answer}’’
‘‘response’’: ‘‘{response}’’
Output:
Figure 16: The prompt used to evaluate the consistency between the reference answers generated by DeepSeek-R1-0528 and the CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8. The placeholders {question}, {reference_answer}, and {response} are replaced with the synthesized question, the reference answer, and the CoT responses, respectively.
Prompt for Discipline Label Evaluation You are reviewing exam questions and their author-provided discipline labels. Judge whether the discipline label is appropriate for the question. Answer ‘Yes’ when the label is plausible and answer ‘No’ when there is a clear mismatch. Output only ‘Yes’ or ‘No’.
‘‘question’’: ‘‘{question}’’
‘‘discipline label’’: ‘‘{discipline}’’
Output:
Figure 17: The prompt used to evaluate the appropriateness of the discipline label. The placeholders {question} and {discipline} are replaced with the specific question and its discipline label.
Prompt for Difficulty Label Evaluation You are reviewing exam questions and their author-provided difficulty labels. Judge whether the difficulty label is appropriate for solving the question. Answer ‘Yes’ when the label is plausible and ‘No’ when there is a clear mismatch. Output only ‘Yes’ or ‘No’.
‘‘question’’: ‘‘{question}’’
‘‘difficulty label’’: ‘‘{difficulty}’’
Output:
Figure 18: The prompt used to evaluate the appropriateness of the difficulty label. The placeholders {question} and {difficulty} are replaced with the specific question and its difficulty label.
Prompt for Type Label Evaluation You are reviewing exam questions and their author-provided type labels. Judge whether the type label is appropriate for the question. Answer ‘Yes’ when the label is plausible and answer ‘No’ when there is a clear mismatch. Output only ‘Yes’ or ‘No’.
‘‘question’’: ‘‘{question}’’
‘‘type label’’: ‘‘{qtype}’’
Output:
Figure 19: The prompt used to evaluate the appropriateness of the question type label. The placeholders {question} and {qtype} are replaced with the specific question and its type label.
Appendix K Design Logic Examples